# 6th Workshop on Patent Text Mining and Semantic Technologies (PatentSemTech) 2025

### v2

@@ Abstract
We introduce PEDANTIC (Patent Definiteness Examination Corpus), a novel dataset of 14k US patent claims from patent applications relating to Natural Language Processing (NLP), annotated with reasons for indefiniteness. We construct PEDANTIC using a fully automatic pipeline that retrieves office action documents from the USPTO and uses Large Language Models (LLMs) to extract the reasons for indefiniteness. A human validation study confirms the pipeline’s accuracy in generating high-quality annotations. To gain insight beyond binary classification metrics, we implement an LLM-as-Judge evaluation that compares the free-form reasoning of every model-cited reason with every examiner-cited reason. We show that LLM agents based on Qwen 2.5 32B and 72B struggle to outperform logistic regression baselines on definiteness prediction, even though they often correctly identify the underlying reasons. PEDANTIC provides a valuable resource for patent AI researchers, enabling the development of advanced examination models. We will publicly release the dataset and code.
@@ 1 Introduction
The patent system plays a crucial role in fostering innovation by granting inventors exclusive rights to their inventions. Central to this system are patent claims, which are concise and precise statements that define the metes and bounds of the protected invention. The process of obtaining a patent involves rigorous examination by patent offices to ensure that the application meets specific criteria, including novelty, non-obviousness, and, critically, definiteness. The latter requires that every claim is sufficiently clear and unambiguous to enable a person skilled in the art (called the Person of Ordinary Skill in the Art, or POSITA) to understand the scope of the invention. In American patent law, this is defined in 35 U.S.C. § 112(b) (comparable to clarity in the EU [^1] https://www.epo.org/en/legal/guidelines-epc/2024/f_iv_4.html), which states that the specification shall conclude with one or more claims particularly pointing out and distinctly claiming the subject matter which the applicant regards as his invention. The Manual of Patent Examining Procedure (MPEP) [1] provides detailed instructions for the examination. We present the most common categories of indefiniteness in Table 1.
@@
Indefiniteness is one of the most common reasons for rejection [5, 6, 7], yet its automatic examination has received very little attention in prior work compared to novelty and non-obviousness. To bridge this gap, we specifically target the automatic examination of definiteness. We argue that predicting definiteness as a binary label is insufficient for practical applications. For patent drafters and examiners, knowing that a claim is indefinite does not provide actionable information. Rather, understanding why it is indefinite is crucial, as it enables targeted improvements to the claim. For this purpose, we introduce PEDANTIC, a dataset of patent claims annotated with detailed justifications for rejection due to indefiniteness. In addition to the claims, PEDANTIC includes fine-grained indefiniteness categories (such as ’antecedent basis’) along with a free-form reasoning, and the affected ranges of the claim for every indefiniteness reason. We create it fully automatically, leveraging Large Language Models (LLMs) to parse USPTO office actions into a structured format that includes these annotations. In total, PEDANTIC includes 14k claims from 3k utility patent applications relating to Natural Language Processing (NLP) filed after 2014. It facilitates a nuanced evaluation, distinguishing between systems that identify indefiniteness based on superficial clues and those that are able to pinpoint the correct underlying issue. In addition to established classification metrics, we implement an LLM-as-Judge [8] evaluation that compares each model-cited rejection reason with every examiner-cited rejection reason. We release our dataset and code publicly to facilitate further research.
@@ 3 PEDANTIC Dataset
In this section, we describe steps taken to create our dataset, including the retrieval of seed patent applications, the retrieval of rejection notices, document parsing, and dataset splitting.
@@ 3.1 Seed Patent Applications
Our dataset creation pipeline first requires a set of seed patent applications. We focus on patent applications in the area of NLP, i.e., applications with the CPC class G06F40 (“Handling natural language data”), but our pipeline is agnostic to this and can be readily used with seed applications from any other field. We query the USPTO Open Data Portal (ODP) [^2] https://data.uspto.gov/ API for applications from 2014 to date with at least one filed rejection notice. We do not consider applications prior to 2014 because the requirements for definiteness changed significantly after the US Supreme Court’s ruling on Nautilus vs. Biosig [26]. Before 2014, a claim was only rejected for indefiniteness if it was “insolubly ambiguous”, whereas afterward, indefiniteness was interpreted as a claim “[failing] to inform, with reasonable certainty, those skilled in the art about the scope of the invention”. Ashtor [21] finds empirical evidence that this ruling has significantly increased the fraction of claims being rejected for indefiniteness and the average claim clarity. To avoid biases that could be exploited by models, we choose to filter out applications filed before 2014.
@@ 3.3 Office Action Parsing
Next, we parse the office actions from Markdown into a structured representation using Gemma 3 27B [27]. First, we select the sections related to indefiniteness by filtering for sections with headings containing “112”. Next, we prompt the LLM with the selected sections and instruct it to extract indefiniteness reasons in a JSON schema. In particular, each rejection contains the text snippet arguing why the claim is indefinite, a category from Table 1, and a list of recited phrases. We instruct the LLM to extract the free-form argumentation and recited phrases verbatim and not to paraphrase, extend or explain them. The prompt is attached in the Appendix A. Lastly, we use fuzzy matching to find the occurrences of recited phrases in the claim text. An example from our dataset is shown in Figure 1. A claim can have multiple rejection reasons, but most have only one (see Section 3.7).
@@ 3.4 Sampling Definite Claims
To ensure robust evaluation, we balance the dataset between definite and indefinite claims. We use claims from applications whose office actions do not contain the term “112(b)”, i.e., applications whose claims are all definite. This makes sure that all claims labelled as definite are indeed not rejected for indefiniteness; using claims from the same applications as the previously extracted indefinite claims could introduce noise if the parsing pipeline does not detect all rejections. To also balance the number of applications in each class, we first compute the average number of indefinite claims per application included in the dataset. We iterate through the definite applications in random order and sample the same number of claims until there are as many definite as indefinite claims. While sampling claims from the applications, we round the number of sampled claim down or up depending on whether there are currently more or less claims per application than in the indefinite samples.
@@ 4.2 LLM Agent
Rather than predicting the binary label (definite/indefinite), we instruct the LLM to use a verbalized expression of likelihood among a set of possible options ranging from “almost no chance” to “almost certain”. Such a confidence score, if reliable, should allow users to configure a level of strictness for issue detection. Each expression is converted to a numerical probability using a lookup table adapted from the empirically determined human perception of probability expressions by Fagen-Ulmschneider [28]. We choose this approach because Tian et al. [29] and Xiong et al. [30] show that asking the LLM for a confidence value is equally or more reliable than logit-based approaches.
@@ 5 Evaluation Metrics
In this section, we propose a set of metrics to evaluate claim indefiniteness prediction models. First, since the core task is binary classification, we use established metrics that measure how well the system can predict the binary label “definite” or “indefinite”. While this binary label is already interesting by itself, practical applications demand a more fine-grained and explainable approach. It is equally or even more important to know why a claim is indefinite to give actionable feedback to drafters. To that end, we implement an LLM-as-Judge [8] approach that determines whether an examiner-cited reason and a model-cited reason point to the same essential issue in the claim. Lastly, we evaluate the task as a multi-label classification where the overlap in the categories of cited reasons also hint towards whether a model decided to accept or reject a claim for the right reasons.
@@ 5.2 Pairwise Reasoning Judge
To assess the quality of the model’s reasoning for indefiniteness, we employ an LLM-as-Judge approach [8]. We prompt Gemma 3 27B to evaluate whether an examiner-cited reason and a model-cited reason point to the same essential issue in the claim on a scale from 1 (worst) to 5 (best) (we normalize the final scores to the range $[0,100]$). The few-shot prompt is attached in Appendix C. To allow the model to analyze the ground truth and the predicted reason before settling on a grade, we first prompt it to find similarities and differences, then ask for a numerical grade in a follow-up message. Following latest research in LLM-as-Judge systems [31, 32], we use the probability-weighted mean as the similarity score of a reason-pair:
@@ 5.3 Multi-Label Classification
We also assess the system as a multi-label classification task, where each label serves as a binary indicator denoting the presence or absence of a specific category within the list of indefiniteness reasons. A high score means that the model has classified a claim for reasons belonging to the same or overlapping indefiniteness categories as the examiner, and thus also represents a kind of explanation for the binary classification. We report the micro and macro average of the category-wise F1 scores. As with the binary classification, we compute the metrics for balanced predictions by first determining the confidence threshold per label under which the fraction of positive samples from the validation set matches.
@@ 6 Results
Figure 4: Distributions of claim-level indefiniteness probability estimates for different binary classifiers and the relation of estimated probability and accuracy. The accuracy per probability bin is plotted as red dots. For bins with less than 100 samples, the dot is faded. The logistic regression model shown here uses the full feature set.
@@
Among the three proposed variants of logistic regression, the variant with linguistic features and the variant with all features perform comparably, while the variant using only TF-IDF features performs substantially worse. This indicates that the handcrafted features are indeed helpful to differentiate definite and indefinite claims, and that TF-IDF features provide only limited additional information.
We show the feature importance values in Figure 3. The most important feature indicating definiteness is the IOU of words between claim and description, i.e., samples with a high word overlap between claim and description are less likely to be rejected for indefiniteness. The highest-weighted features indicating indefiniteness are the number of unique word stems, the type token ratio and the number of stopwords. Among the binary features showing whether a trigger word is contained in the claim, “step for” is the most indicative of indefiniteness. The flag whether the claim is independent indicates definiteness, which is in line with our dataset statistics, albeit with moderate importance. Many of the text complexity and readability metrics have high importance scores. A higher text complexity is associated with definiteness (with Gunning Fog, Automated Readability Index, and Flesch Kincaid Grade having a negative feature weight) and better readability is associated with indefiniteness (with Flesch Reading Ease and Dale Chall Readability Score having a positive feature weight). While this might seem counterintuitive when relating definiteness to clarity or understandability, it is intuitive when relating definiteness to the absence ambiguity, as removing ambiguities often requires additional specification, making the text more complex.
@@
Figure 5: Distributions of reason-level confidence scores and the relation of confidence and correctness according to LLM-as-Judge. The reported judge scores are the soft micro-averaged precision scores, and thus represent the average max similarity per model-cited reason. Dots with less than 100 samples are faded. The dashed line is a linear regression that ignores the faded points. $\rho$ corresponds to the pearson correlation between confidence and judge score.
@@
LLM agents and logistic regression achieve only moderate F1 scores. All models achieve a much better micro F1 than macro F1, indicating that the performance varies between categories. We find that all models perform much better on the well-represented categories than on rare ones. Given the substantial noise in the labels for the category undefined term (see Section 3.6), additional investigations are necessary to draw further conclusions.
@@
Our LLM-as-Judge evaluation provides further insight into LLM agents’ performance by directly analyzing the quality of their identified reasons for indefiniteness. Table 5 shows the results for Qwen 2.5 32B and 72B. Interestingly, the smaller model clearly outperforms the larger one in all judge metrics, despite exhibiting slightly lower binary classification performance. There seems to be a disconnect between the models’ ability to identify potential reasons for indefiniteness and its ability to accurately decide whether a claim is indefinite. The 32B model is able to correctly identify 35.4% of the examiner-cited reasons (micro $R_{\geq 75}$, 25.0% for 72B). The precision is much lower than the recall, likely due to the large number of low-confidence reasons (see Figure 5). The models tend to list many reasons, even if they consider them unlikely to cause indefiniteness.
@@ 7 Discussion
In this section, we provide qualitative insight into the reasons for indefiniteness generated by Qwen 2.5 32B compared to those written by the human examiner. Table 6 lists three examples. In Example 1, the LLM agent correctly identified the reason for indefiniteness, albeit with a less verbose explanation. In Example 2, the LLM agent also found the correct underlying issue. However, its explanation is not equally helpful as the one provided by the human examiner, which additionally suggests that the claims might accidentally depend from the wrong base claim. In Example 3, the model-generated reason points to similar problematic phrases, but without identifying the underlying issue. The LLM-as-Judge correctly identifies that there is little substantial overlap and assigns a lower score of 25%. Generally, the model-generated reasons seem to be less detailed, and often lack the deep analysis and reasoning performed by human examiners.

### v3

@@ Abstract
We introduce PEDANTIC (Patent Definiteness Examination Corpus), a novel dataset of 14k US patent claims from patent applications relating to Natural Language Processing (NLP), annotated with reasons for indefiniteness. We construct PEDANTIC using a fully automatic pipeline that retrieves office action documents from the USPTO and uses Large Language Models (LLMs) to extract the reasons for indefiniteness. A human validation study confirms the pipeline’s accuracy in generating high-quality annotations. To gain insight beyond binary classification metrics, we implement an LLM-as-Judge evaluation that compares the free-form reasoning of every model-cited reason with every examiner-cited reason. We show that LLM agents based on Qwen 2.5 32B and 72B struggle to outperform logistic regression baselines on definiteness prediction, even though they often correctly identify the underlying reasons. PEDANTIC provides a valuable resource for patent AI researchers, enabling the development of advanced examination models. We release the dataset and code at https://github.com/boschresearch/pedantic-patentsemtech.
@@ 1 Introduction
The patent system plays a crucial role in fostering innovation by granting inventors exclusive rights to their inventions. Central to this system are patent claims, which are concise and precise statements that define the metes and bounds of the protected invention. The process of obtaining a patent involves rigorous examination by patent offices to ensure that the application meets specific criteria, including novelty, non-obviousness, and, critically, definiteness. The latter requires that every claim is sufficiently clear and unambiguous to enable a person skilled in the art (called the Person of Ordinary Skill in the Art, or POSITA) to understand the scope of the invention. In American patent law, this is defined in 35 U.S.C. § 112(b) (comparable to clarity in the EU [^1] https://www.epo.org/en/legal/epc/2020/a84.html
https://www.epo.org/en/legal/guidelines-epc/2024/f_iv_4.html
), which states that the specification shall conclude with one or more claims particularly pointing out and distinctly claiming the subject matter which the applicant regards as his invention. The Manual of Patent Examining Procedure (MPEP) [1] provides detailed instructions for the examination. We present the most common categories of indefiniteness in Table 1.
@@
Indefiniteness is one of the most common reasons for rejection [5, 6, 7], yet its automatic examination has received very little attention in prior work compared to novelty and non-obviousness. To bridge this gap, we specifically target the automatic examination of definiteness. We argue that predicting definiteness as a binary label is insufficient for practical applications. For patent drafters and examiners, knowing that a claim is indefinite does not provide actionable information. Rather, understanding why it is indefinite is crucial, as it enables targeted improvements to the claim. For this purpose, we introduce PEDANTIC, a dataset of patent claims annotated with detailed justifications for rejection due to indefiniteness. In addition to the claims, PEDANTIC includes fine-grained indefiniteness categories (such as ’antecedent basis’) along with a free-form reasoning, and the affected ranges of the claim for every indefiniteness reason. We create it fully automatically, leveraging Large Language Models (LLMs) to parse USPTO office actions into a structured format that includes these annotations. In total, PEDANTIC includes 14k claims from 3k utility patent applications relating to Natural Language Processing (NLP) filed after 2014. It facilitates a nuanced evaluation, distinguishing between systems that identify indefiniteness based on superficial clues and those that are able to pinpoint the correct underlying issue. In addition to established classification metrics, we implement a reference-based LLM-as-Judge [8] evaluation that compares each model-cited rejection reason with every examiner-cited rejection reason. We release our dataset and code publicly to facilitate further research.
@@ 3 PEDANTIC Dataset
In this section, we describe the creation of our dataset, including the retrieval of seed patent applications, the retrieval of rejection notices, document parsing, and dataset splitting.
@@ 3.1 Seed Patent Applications
Our dataset creation pipeline first requires a set of seed patent applications. We focus on patent applications in the area of NLP, i.e., applications with the CPC class G06F40 (“Handling natural language data”), but our pipeline is agnostic to this and can be readily used with seed applications from any other field. We query the USPTO Open Data Portal (ODP) [^2] https://data.uspto.gov/ API for applications from 2014 to date with at least one filed rejection notice. We do not consider applications prior to 2014 because the requirements for definiteness changed significantly after the US Supreme Court’s ruling on Nautilus vs. Biosig [26]. Before 2014, a claim was only rejected for indefiniteness if it was “insolubly ambiguous”, whereas afterward, indefiniteness was interpreted as a claim “[failing] to inform, with reasonable certainty, those skilled in the art about the scope of the invention”. Ashtor [21] finds empirical evidence that this ruling has significantly increased the fraction of claims being rejected for indefiniteness and the average claim clarity. To avoid exploitable biases, we filter out applications filed before 2014.
@@ 3.3 Office Action Parsing
We parse the office actions’ full-text into a structured representation using Gemma 3 27B [27]. First, we select the sections related to indefiniteness by filtering for sections with headings containing “112”. Next, we prompt the LLM with the selected sections and instruct it to extract indefiniteness reasons in a JSON schema. In particular, each rejection contains the text snippet arguing why the claim is indefinite, a category from Table 1, and a list of recited phrases. We instruct the LLM to extract the free-form argumentation and recited phrases verbatim and not to paraphrase, extend or explain them. The prompt is attached in Appendix A. Lastly, we use fuzzy matching to find the occurrences of recited phrases in the claim text. An example from our dataset is shown in Figure 1. A claim can have multiple rejection reasons, but most have only one (see Section 3.7).
@@ 3.4 Sampling Definite Claims
To ensure robust evaluation, we balance the dataset between definite and indefinite claims. We use claims from applications whose office actions do not contain the term “112(b)”, i.e., applications whose claims are all definite. This makes sure that all claims labelled as definite are indeed not rejected for indefiniteness; using claims from the same applications as the previously extracted indefinite claims could introduce noise if the parsing pipeline does not detect all rejections. To also balance the number of applications in each class, we first compute the average number of indefinite claims per application included in the dataset. We iterate through the definite applications in random order and sample the same number of claims until there are as many definite as indefinite claims. While sampling claims from the applications, we round the number of sampled claims down or up depending on whether there are currently more or less claims per application than in the indefinite samples.
@@ 4.2 LLM Agent
Rather than generating the binary label directly (definite/indefinite), we instruct the LLM to use a verbalized expression of likelihood among a set of possible options ranging from “almost no chance” to “almost certain”. Confidence scores, if reliable, should allow users to configure a level of strictness for issue detection. Each expression is converted to a numerical probability according to the empirically determined human perception of probability expressions by Fagen-Ulmschneider [28]. We choose this approach because Tian et al. [29] and Xiong et al. [30] show that asking the LLM for a confidence value is equally or more reliable than logit-based confidence estimation.
@@ 5 Evaluation Metrics
In this section, we propose a set of metrics to evaluate claim indefiniteness prediction models. First, since the core task is binary classification, we use established metrics that measure how well the system can predict the binary label “definite” or “indefinite”. While this binary label is already interesting by itself, practical applications demand a more fine-grained and explainable approach. It is equally or even more important to know why a claim is indefinite to give actionable feedback to drafters. To that end, we implement an LLM-as-Judge [8] approach that determines whether an examiner-cited reason and a model-cited reason point to the same essential issue in the claim. Lastly, we evaluate the task as a multi-label classification, where the overlap in the categories of cited reasons also indicate whether a model decided to accept or reject a claim for the right reasons.
@@ 5.2 Pairwise Reasoning Judge
To assess the quality of the model’s reasoning for indefiniteness, we employ a reference-based LLM-as-Judge approach [8]. We prompt Gemma 3 27B to evaluate whether an examiner-cited reason and a model-cited reason point to the same essential issue in the claim on a scale from 1 (worst) to 5 (best) (we normalize the final scores to the range $[0,100]$). The few-shot prompt is attached in Appendix C. To allow the model to analyze the ground truth and the predicted reason before settling on a grade, we first prompt it to find similarities and differences, then ask for a numerical grade in a follow-up message. Following latest research in LLM-as-Judge systems [31, 32], we use the probability-weighted mean as the similarity score of a reason-pair:
@@ 5.3 Multi-Label Classification
We also assess the system as a multi-label classification task, where each label serves as a binary indicator denoting the presence or absence of a specific category within the list of indefiniteness reasons. A high score means that the model has classified a claim for reasons belonging to the same or overlapping indefiniteness categories as the examiner, and thus also represents a kind of explanation for the binary classification. We report the micro and macro average of the category-wise F1 scores. As with the binary classification, we balance the predictions using the confidence threshold per category under which the fraction of positive predictions matches the distribution in the validation set.
@@ 6 Results
Figure 4: Distributions of claim-level indefiniteness probability estimates for different binary classifiers and the relation of estimated probability and accuracy. The accuracy per probability bin is plotted as red dots. For bins with less than 1% of samples, the dot is faded. The logistic regression model shown here uses the full feature set.
@@
Among the three proposed variants of logistic regression, the one with linguistic features and the one with all features perform comparably, while using only TF-IDF features performs substantially worse. This indicates that the handcrafted features are indeed helpful to differentiate definite and indefinite claims, and that TF-IDF features provide only limited additional information.
We show the feature importance values in Figure 3. The most important feature indicating definiteness is the IOU of words between claim and description, i.e., samples with a high word overlap between claim and description are less likely to be rejected for indefiniteness. The highest-weighted features indicating indefiniteness are the number of unique word stems, the type token ratio and the number of stopwords. Among the binary features showing whether a trigger word is contained in the claim, “step for” is the most indicative of indefiniteness. The flag whether the claim is independent indicates definiteness, which is in line with our dataset statistics, albeit with moderate importance. Many of the text complexity and readability metrics have high importance scores. A higher text complexity is associated with definiteness (with Gunning Fog, Automated Readability Index, and Flesch Kincaid Grade having a negative feature weight) and better readability is associated with indefiniteness (with Flesch Reading Ease and Dale Chall Readability Score having a positive feature weight). While this might seem counterintuitive when relating definiteness to clarity or understandability, it makes sense when considering definiteness as the absence of ambiguity, as removing ambiguities often adds complexity through additional specification.
@@
Figure 5: Distributions of reason-level confidence scores and the relation of confidence and correctness according to LLM-as-Judge. The reported judge scores are the soft micro-averaged precision scores, and thus represent the average max similarity per model-cited reason. For bins with less than 1% of samples, the dot is faded. The dashed line is a linear regression that ignores the faded points. $\rho$ corresponds to the pearson correlation between confidence and judge score.
@@
LLM agents and logistic regression achieve only moderate F1 scores. All models achieve a much better micro F1 than macro F1, indicating that the performance varies between categories. All models perform much better on the well-represented categories than on rare ones. Given the substantial noise in the labels for the category undefined term (see Section 3.6), additional investigations are necessary to draw further conclusions.
@@
Our LLM-as-Judge evaluation provides further insight into LLM agents’ performance by directly analyzing the quality of their identified reasons for indefiniteness. Table 5 shows the results for Qwen 2.5 32B and 72B. Interestingly, the smaller model clearly outperforms the larger one in all judge metrics, despite exhibiting slightly lower binary classification performance. There seems to be a disconnect between the models’ ability to identify potential reasons for indefiniteness and its ability to accurately decide whether a claim is indefinite. The 32B model is able to correctly identify 35.4% of the examiner-cited reasons (micro $R_{\geq 75}$, 25.0% for 72B). Both models list many reasons, even if they consider them unlikely to cause indefiniteness, causing the precision to be lower than the recall.
@@ 7 Discussion
We here compare example reasons for indefiniteness generated by Qwen 2.5 32B with those written by the human examiner. Table 6 lists three examples. In Example 1, the LLM agent correctly identified the reason for indefiniteness, albeit with a less verbose explanation. In Example 2, the LLM agent also found the correct underlying issue. However, its explanation is not equally helpful as the one provided by the human examiner, which additionally suggests that the claims might accidentally depend from the wrong base claim. In Example 3, the model-generated reason points to similar problematic phrases, but without identifying the underlying issue. The LLM-as-Judge correctly identifies that there is little substantial overlap and assigns a lower score of 25%. Generally, the model-generated reasons seem to be less detailed, and often lack the deep analysis and reasoning performed by human examiners.
While our dataset creation process is domain-agnostic, our study focuses on patents from the NLP domain. Therefore, the claim language and terminology is relatively homogeneous. Other domains might differ in terms of terminology, distribution of claim types and categories of indefiniteness, and legal conventions. Future work should generalize our study design and verify that our findings transfer to other domains.
@@ 8 Conclusion and Outlook
Acknowledgements.
We would like to thank the patent attorneys Philipp Mangold and Charlotte Hellmann for insightful discussions and helpful pointers.